OpenStudy (anonymous):

I'm trying to determine what the perimeter is when the following are given: 5b^2+8b-12 b-4 18b^2-3b+11

OpenStudy (mathstudent55):

Is it a triangle?

OpenStudy (mathstudent55):

Are the expressions the lengths of the sides? The perimeter of a polygon is the sum of the measures of the sides.

OpenStudy (anonymous):

yes, it is a triangle so the b-4 is the base

OpenStudy (anonymous):

I added the equations: 23b^2+5b-5

OpenStudy (mathstudent55):

You need to add the lengths of the three sides: \((5b^2+8b-12)+ (b-4) +(18b^2-3b+11)\) The parentheses can be dropped because they are unnecessary since you adding the three expressions: \(=5b^2+8b-12+ b-4 +18b^2-3b+11\) You need to combine like terms. Like terms have the same variable parts. \(=5b^2+18b^2+8b+ b-3b-12-4 +11\) \(=23b^2+6b-5\) You were pretty close. You just missed the b term of the second side. Keep in mind that \(b = 1b\), so \(8b+ b-3b\) \( = 8b + 1b - 3b\) \( = 9b - 3b\) \( = 6b\)